Hi, I have a problem that reads: "The balloon traveled 48 miles in 4 hours. It's speed was then increased by 8 mph. If the total they traveled was 108 miles, how much time did it need for the entire trip?" This is what I did: I subtracted 48 miles from the total 108 miles it traveled and got 60. Then I divided 48 miles by 4 hours and got 12 mph. Then I added 8 mph increase to the 12 mph and got 20 mph, which I used to divide the 60 miles by, and got 30 hours. Is that right? (the answer is a total of 34 hours?) Also, if this is for Algebra, can I just show my work as simple subtraction, addition, and dividing or do I have to write some equation? It doesn't say to do that.. Thank you!
Answer by Nate(3500)

You are right except for the division:
60/20 = 3 hours
7 hours total
3 hours (20 mph) + 4 hours (12 mph) = 108 ?
60 + 48 = 108 ?
True